OpenStudy (anonymous):

write an equation for each line. slopes = 3/4 ; (-3,5)

OpenStudy (anonymous):

Use the equation for a line y-y1=m(x-x1)

OpenStudy (anonymous):

i did but after the y-5=3/4x i got stuck

OpenStudy (anonymous):

\[y-5=\frac{ 3 }{ 4 }(x-(-3))\] \[y-5=\frac{ 3 }{ 4 }(x+3)\] Do you follow?

OpenStudy (anonymous):

yes but how do you distribute 3/4 to 3

OpenStudy (anonymous):

What do you mean distribute? We just plug in the x and y coordinates and gradient into the equation.